OpenStudy (anonymous):

Write the prime factorization of each number. Use division. Show your work. For number 711.

OpenStudy (anonymous):

do you know instantly that this number is divisible by 9?

OpenStudy (anonymous):

3^2×79

OpenStudy (anonymous):

3*3*79

OpenStudy (anonymous):

no i didn't , i didnt think it was divisable by anything . :/

OpenStudy (anonymous):

\[7+1+1=9\] so it is divisible by 9. add up the digits

OpenStudy (anonymous):

7+1+1 = 9, so it is divisible by 9, also divisible by 3

OpenStudy (anonymous):

then divide 711 by 9 you will get \[711\div9=79\] so \[711=9\times 79\] of course 9 is not prime , its \[3\times 3\] so your answer is \[3\times 3\times 79\] or \[3^2\times 79\]